OpenStudy (anonymous):

If 0 < x < 8 and -1 < y < 3, whichofthefollowinggivesthesetofallpossible values of xy ? (A) xy = 4 (B) 0 < xy < 24 (C) -1 < xy < 11 (D)-1

OpenStudy (anonymous):

max value of xy=8*3=24 min value of xy=-1*8=-8

OpenStudy (anonymous):

E

OpenStudy (anonymous):

thanks